# Bayesian neural networks with interpretable priors from Mercer kernels

### v1

@@ 4.1 Dependence of the prior on the eigenvalues and eigenfunctions
The results of the two-sample KS test are presented in Fig. 7. For the test, set the significance level to be $\alpha=0.05$. The KS statistic suggests that the marginal distributions of the BNN match that of Brownian motion when approximately $1,000$ eigenvalues/eigenfunctions are accounted for. We feel uncomfortable making this claim for $K=500$, as there is a region where the KS statistic largely falls above $D_{0.05}$ (for $t\in(0.2,0.4)$). The same holds for the $p$-values. What is notable is that in all cases, the KS statistic lies outside the acceptable threshold when $t$ is close to 0. We attribute this to the specific parameterization we have chosen, i.e., $u_{\theta}(t)=tf(t;\theta)$, which could be forcing the BNN to remain too small close to 0. Again, we made this choice so that the initial condition $u_{\theta}(0)=0$ is satisfied.
@@ 4.2 Dependence of the prior on the network width
The results of these numerical experiments show that the width of the neural network plays a key role in how well the BNN prior approximates the target GP. As the network becomes wider, the approximation becomes higher-fidelity as evidenced in the empirical covariance matrices. However, the KS test suggests that the improvement plateaus, as both the $1,000$ neuron and $2,500$ neuron network yield similar results. This suggests diminishing returns once the network is expressive enough relative to the number of eigenfunctions used. The use of the a wider network will reduce the error, but this comes at an increased computational cost. Together with the results of Sec. 4.1, this highlights that both the spectral truncation $K$ and the network width $N_{1}$ jointly determine the of the approximation, with convergence expected in the limit $K,N_{1}\to\infty$.
@@ 5 Example applications
We provide some example applications of the Mercer prior in practice. We highlight situations in which it is beneficial to replace a GP with a BNN, while also showing that the addition of a Mercer prior improves the performance of a BNN when compared to an i.i.d. Gaussian prior. The examples covered all contain real-world data, and not simple synthetic examples, which we hope helps makes the case that the Mercer prior is worth the effort to implement in practice.
@@ 5.1 Gaussian process regression with heteroscedastic noise
For the first example, we highlight the ease with which the networks can be deployed in GP regression tasks with difficult posteriors. This example demonstrates how the method allows for easy hierarchical GP regression. We consider the dataset of [45], which consists of measurements of the acceleration of a helmet during a motorcycle cycle at various points in time. Note that we standardize the measurements and rescale the domain to $[0,1]$. It is well-known that these individual measurements are subject to heteroscedastic noise. We build a hierarchical Bayesian model to perform inference on this dataset, replacing any GPs with BNNs sampled from a Mercer prior.
@@
where $m$ and $\sigma^{2}$ are functions which represent the evolving mean acceleration and heteroscedastic variance of the measurement, respectively. We represent both as centered GPs $m\sim\mathcal{GP}(0,\Delta^{-2})$ and $\sigma^{2}\sim\mathcal{GP}(0,\Delta^{-2})$, which places the inductive bias that the mean and variance should should live in the space $H^{1}([0,1])$ at the bare minimum, by Lemma 3.2.
@@
We replace each GP with a BNN equipped with the corresponding Mercer prior, which we denote by $m_{\theta}$ and $\sigma^{2}_{\psi}$. So, each BNN is governed by an individual Mercer prior, i.e., $p(\theta)\propto\exp\left(-1/2\langle m_{\theta},\Delta^{2}m_{\theta}\rangle\right)$ and $p(\psi)\propto\exp\left(-1/2\langle\sigma^{2}_{\psi},\Delta^{2}\sigma^{2}_{\psi}\rangle\right)$ This gives the hierarchical Bayesian model
@@
which is then sampled with SGLD. The resulting posterior is shown in Fig. 13. We see that the BNNs are both learn the overall trend of the data and identify the heteroscedastic noise present. Also, we generate the posterior on a super-fine mesh of one million points to further emphasize the ease at which the method scales.
@@ 5.2 A periodic BNN
(a) BNN posterior with the Mercer prior
(b) BNN posterior with an i.i.d. Gaussian prior
Figure 14: Posterior predictive distribution on the $\mathrm{CO}_{2}$ for a BNN with a Mercer prior and a BNN with an i.i.d. Gaussian prior. The shaded region represents the $95\%$ credible interval. Samples from the posterior are shown in green.
@@ 5.3 Nonlinear Bayesian inverse problem
For the final example, we demonstrate how a BNN with a Mercer prior can serve as a replacement for a Gaussian measure in a nonlinear Bayesian inverse problem. This example is related to the design of spacecraft thermal protection systems (TPS), where it is critical to understand the insulation properties of certain materials so that a TPS may be appropriately sized. Real world examples include the Apollo spacecraft [40] or the Space Shuttle [11]. For this example, we consider the dataset found in [13], where the thermal conductivity $\kappa$ of the Opacified Fibrous Insulation (OFI) is identified through experiments. The dataset consists of steady-state temperature measurements of the OFI ranging from $300$ K to $1900$ K. The goal is to identify the thermal conductivity of the OFI from the temperature measurements with uncertainty, which we then compare to the reported values.
@@
where the domain is scaled from $0\leq y\leq 1$, $T$ denotes the temperature, and $\kappa$ denotes the thermal conductivity of the OFI, which is known to depend on the temperature. The boundary conditions are $T(0)=300$ and $T(1)=1900$. The resulting temperature measurements are denoted by $\mathbf{d}=(T_{i})_{i=1}^{n}$ and are assumed to be corrupted with zero-mean Gaussian noise with a standard deviation of $\sigma=2.5$ K. This results in an infinite-dimensional nonlinear inverse problem.
Under the classical approach, one would start by assigning a Gaussian measure to the thermal conductivity $\mu_{0}=\mathcal{N}(m,S)$, with $\log\kappa\sim\mu_{0}$ so that Bayes’s theorem can be used to derive the posterior. The measurement operator $R$ is given by the solution to eqn. (5.2) at the measurement locations. Therefore the posterior $\mu^{\mathbf{d}}$ is given by the Radon-Nikodym derivative
@@
We demonstrate how this inverse problem can be solved by replacing the prior Gaussian measure with a BNN with the equivalent Mercer prior. For the prior knowledge, we need to impose the condition that $\kappa$ has at least one derivative, as $\partial\kappa/\partial T$ appears in eq. (5.2). Rather than defining the prior in the $\log$-space, we instead apply a softplus transformation, $\mathrm{softplus}(x)=\ln(1+e^{x})$, which helps prevent underflow. Putting this together, we select the prior $\mathrm{softplus}(\kappa)\sim\mathcal{N}(0,\Delta^{-2})$. We represent this prior as a BNN $\kappa_{\theta}$ with corresponding Mercer prior, with the eigenvalues and eigenfunctions being as discussed in § 3.2.
@@
Before sampling from the posterior, we initialize the parameters at the maximum a posteriori estimate $\theta_{0}=\arg\max p(\theta|\mathbf{d})$. The resulting BNN posterior is presented in Fig. 15, where we find that the posterior accurately reconstructs the latent thermal conductivity, although there is some discrepancy at higher temperatures. As this is an example with real experimental data, some amount of error is to be expected.
@@ 6 Conclusions
In this work, we introduced Mercer priors, a new class of priors for BNNs derived from the Mercer representation of a covariance kernel. The prior enforces a covariance structure on the parameters of the BNN such that in the output space the networks behave similarly to the corresponding GP with the chosen covariance. This allows for BNNs to inherit the interpretability of GP priors while retaining the scalability and flexibility of neural networks. A challenge that remains is the treatment of any hyperparameters appearing in the GP covariance, which is discussed in detail in Appendix A.
@@ Acknowledgments
Appendix A Proposed directions for hyperparameter selection
which defines the properly normalized prior $p(\theta|\lambda)=Z(\lambda)^{-1}\exp(-E(\theta;\lambda))$. To connect the technique back to its inspiration in field theories, $E$ can be viewed as the energy of the system, and $Z(\lambda)$ is the corresponding partition function. Under this view, the prior assigns higher probability density to neural networks which are closer to minimum energy state.
A fully Bayesian treatment over both $\theta$ and $\lambda$ requires accounting for the variation of $Z(\lambda)$, as the posterior would be
once one picks a hyperprior $p(\lambda)$. Of course, this is unlikely to be tractable due to the complicated map defined by the neural network architecture within $Z(\lambda)$. Ignoring this subtlety and dropping the partition function introduces bias into the inference whenever $\lambda$ is learned jointly with $\theta$. In our examples, $\lambda$ was always fixed so that $Z(\lambda)$ is simply a constant and can be dropped when sampling with SGLD.
To address this, one approach would be to work with the marginal $p(\lambda|y)=\int_{\Theta}p(\theta,\lambda|y)d\theta$, which can be characterized through the following identity
$\nabla_{\lambda}\log p(\lambda|y)=\mathbb{E}_{p(\theta|y,\lambda)}[-\nabla_{\lambda}E(\theta;\lambda)]-\mathbb{E}_{p(\theta|\lambda)}[-\nabla_{\lambda}E(\theta;\lambda)]-\nabla_{\lambda}\log p(\lambda),$ (A.1)
which is derived in [3]. We will work in the simplified case where $p(\lambda)$ is flat, so that only the expectations appear. Then, the practice reduces to model selection by marginal likelihood [59]. This expression states that the gradient of the $\log$-evidence with respect to the hyperparameters is the difference between the expected energy gradients under the posterior and prior, respectively. Intuitively, the first term dictates how the energy landscape shifts under the posterior, ie., it points towards the direction of $\lambda$ which reduces the energy. The second term is a correction coming from the change in normalization constant. It is a baseline shift in energy that happens even without data. Again to connect back to physics, the derivative of the energy $\nabla_{\lambda}E(\theta;\lambda)$ can be viewed as a force, so eq. (A.1) describes the expected excess force that $\lambda$ is subjected to induced by the data. Probabilistically, this is exactly the score function gradient estimator for the likelihood $\nabla_{\lambda}\mathbb{E}_{p(\theta|\lambda)}[p(y|\theta)]=\mathbb{E}_{p(\theta|\lambda)}[p(y|\theta)\nabla_{\lambda}\log p(\theta|\lambda)]$ [60]. This can be seen by first observing that
by using the property of the partition function that $\nabla_{\lambda}\log Z(\lambda)=-\mathbb{E}_{p(\theta|\lambda)}[\nabla_{\lambda}E(\theta;\lambda)]$. Taking the posterior expectation reveals exactly the form of eq. (A.1).
In practice, the identity eq. (A.1) enables empirical Bayes optimization of $\lambda$ using Monte Carlo estimates of the two expectations, which can be obtained via SGLD samples from the posterior and prior with fixed hyperparameters. A fully Bayesian approach would be possible with nested sampling as found in [3]. This approach uses alternating SGLD updates for $\theta$ to approximate the expectations, which then inform and SGLD step for $\lambda$ in order to sample the joint posterior without explicitly computing $Z(\lambda)$. While exact, methods that require evaluating or estimating $Z(\lambda)$, such as pseudo-marginal Metropolis-Hastings, are typically computationally demanding. One could also optimize $\lambda$ using the score function gradient estimator, e.g., with Adam [26].

### v2

@@ 4.1 Dependence of the prior on the eigenvalues and eigenfunctions
The results of the two-sample KS test are presented in Fig. 7. For the test, we set the significance level to be $\alpha=0.05$. The KS statistic suggests that the marginal distributions of the BNN match that of Brownian motion when approximately $1,000$ eigenvalues/eigenfunctions are accounted for. We feel uncomfortable making this claim for $K=500$, as there is a region where the KS statistic largely falls above $D_{0.05}$ (for $t\in(0.2,0.4)$). The same holds for the $p$-values. What is notable is that in all cases, the KS statistic lies outside the acceptable threshold when $t$ is close to 0. We attribute this to the specific parameterization we have chosen, i.e., $u_{\theta}(t)=tf(t;\theta)$, which could be forcing the BNN to remain too small close to 0. Again, we made this choice so that the initial condition $u_{\theta}(0)=0$ is satisfied to ensure that the operation $S^{-1}u_{\theta}$ is well-defined.
@@ 4.2 Dependence of the prior on the network width
The results of these numerical experiments show that the width of the neural network plays a key role in how well the BNN prior approximates the target GP. As the network becomes wider, the approximation becomes higher-fidelity as evidenced in the empirical covariance matrices. However, the KS test suggests that the improvement plateaus, as both the $1,000$ neuron and $2,500$ neuron network yield similar results. This suggests diminishing returns once the network is expressive enough relative to the number of eigenfunctions used. The use of a wider network will reduce the error, but this comes at an increased computational cost. Together with the results of Sec. 4.1, this highlights that both the spectral truncation $K$ and the network width $N_{1}$ jointly determine the quality of the approximation, with convergence expected in the limit $K,N_{1}\to\infty$.
@@ 5 Example applications
We provide some example applications of the Mercer prior in practice. We highlight situations in which it is beneficial to replace a GP with a BNN, while also showing that the addition of a Mercer prior improves the performance of a BNN when compared to an i.i.d. Gaussian prior. The examples covered all contain real-world data, and not simple synthetic examples, which we hope helps makes the case that the Mercer prior is worth the effort to implement in practice. In each example, we report how each prior is chosen, which closely follows the procedure when selecting a GP prior. We also report the hyperparameters used to create the examples, which are chosen following the usual guidelines. The Mercer prior adds an additional layer of complexity, where one must decide how many eigenvalues and eigenfunctions to keep, as well as the subsampling sizes used to generate the unbiased estimate. Compared to the Brownian motion examples, we find much less domain points are required when building an unbiased estimate of the integrals contained in the prior in order to achieve good posterior performance. We conjecture that this is related to the difference in smoothness of the sample paths. Following the empirical evidence of Sec. 4.1, we settle on keeping the first $100$ eigenvalues and eigenfunctions for kernels based on the Laplacian (for two of the examples), while one of the examples uses an engineered kernel and is discussed later. In all of our examples, we subsample $10\%$ of the total eigenvalues and corresponding eigenfunctions at each SGLD step.
@@ 5.1 Gaussian process regression with heteroscedastic noise
For the first example, we highlight the ease with which the networks can be deployed in GP regression tasks with difficult posteriors. This example demonstrates how the method allows for easy hierarchical GP regression. We consider the dataset of [46], which consists of measurements of the acceleration of a helmet during a motorcycle crash at various points in time. Note that we standardize the measurements and rescale the domain to $[0,1]$. It is well-known that these individual measurements are subject to heteroscedastic noise. We build a hierarchical Bayesian model to perform inference on this dataset, replacing any GPs with BNNs sampled from a Mercer prior.
@@
where $m$ and $\sigma^{2}$ are functions which represent the evolving mean acceleration and heteroscedastic variance of the measurement, respectively. We represent both as centered GPs $m\sim\mathcal{GP}(0,\Delta^{-2})$ and $\sigma\sim\mathcal{GP}(0,\Delta^{-2})$, which places the inductive bias that the mean and standard deviation should live in the space $H^{1}([0,1])$ at the bare minimum, by Lemma 3.2. We select a unit variance for both $m$ and $\sigma$ because the data has been standardized.
(a) Mercer prior samples for the mean function
(b) Posterior over the acceleration
@@
We replace each GP with a BNN equipped with the corresponding Mercer prior, which we denote by $m_{\theta}$ and $\sigma_{\psi}$. We use the same network architecture for each, which consists of a Fourier feature input layer with $32$ terms followed by a single layer with $1,500$ neurons. So, each BNN is governed by an individual Mercer prior, i.e., $p(\theta)\propto\exp\left(-1/2\langle m_{\theta},\Delta^{2}m_{\theta}\rangle\right)$ and $p(\psi)\propto\exp\left(-1/2\langle\sigma_{\psi},\Delta^{2}\sigma_{\psi}\rangle\right)$. This gives the hierarchical Bayesian model
@@
which is then sampled with SGLD. When sampling, the unbiased estimate is built by keeping the first $100$ eigenpairs of $\Delta^{-1}$. Then, at each iteration we randomly subsample $10$ eigenpairs, $64$ points in the time domain, and $24$ of the available $94$ datapoints, which represents roughly $25\%$ of the total data. The initial learning rate is set to $0.001$. The resulting posterior is shown in Fig. 13. We see that the BNNs both learn the overall trend of the data and identify the heteroscedastic noise present. Also, we generate the posterior on a super-fine mesh of one million points to further emphasize the ease at which the method scales.
@@ 5.2 A periodic BNN
(a) Samples from Mercer prior with periodic eigenfunctions
(b) BNN posterior with the Mercer prior
(c) BNN posterior with an i.i.d. Gaussian prior
Figure 14: Posterior predictive distribution on the $\mathrm{CO}_{2}$ data for a BNN with a Mercer prior and a BNN with an i.i.d. Gaussian prior. The shaded region represents the $95\%$ credible interval. Samples from the posterior are shown in green.
@@
We use different neural network architectures for each prior to keep the example fair. The BNN equipped with the Mercer prior follows our typical case: a Fourier feature input layer with $32$ terms, followed by a single layer with $1,500$ neurons. For the i.i.d. Gaussian BNN, we use a more standard architecture since wide, shallow neural networks, which we find works well for the Mercer prior, are uncommon in traditional machine learning. Therefore, we select a deep neural network with $3$ hidden layers and $64$ neurons in each layer. In both cases, we push the network parameters to the maximum likelihood estimate (MLE) before sampling with SGLD. This helps decrease the amount of burn-in steps needed for SGLD (recall that the first phase of SGLD acts as a stochastic optimizer before generating samples). The MLE is identified with $10,000$ steps of Adam with an initial learning rate of $0.01$. For sampling/training with Adam, we select a minibatch size of $16$ of the $48$ total datapoints. In this example, we subsample $24$ points in time in the domain and $5$ eigenpairs when building the estimate required for the Mercer prior. The SGLD initial learning rate is $1\times 10^{-5}$, and we take $100,000$ posterior samples for each BNN.
@@ 5.3 Nonlinear Bayesian inverse problem
For the final example, we demonstrate how a BNN with a Mercer prior can serve as a replacement for a Gaussian measure in a nonlinear Bayesian inverse problem. This example is related to the design of spacecraft thermal protection systems (TPS), where it is critical to understand the insulation properties of certain materials so that a TPS may be appropriately sized. Real world examples include the Apollo spacecraft [40] or the Space Shuttle [11]. For this example, we consider the dataset found in [13], where the thermal conductivity $\kappa$ of the Opacified Fibrous Insulation (OFI) is identified through experiments. The dataset consists of steady-state temperature measurements of the OFI ranging from $300$ K to $1900$ K. In total, we have $51$ temperature measurements of the material. The goal is to identify the thermal conductivity of the OFI from the temperature measurements with uncertainty, which we then compare to the published experimental values.
@@
where the domain is scaled from $0\leq y\leq 1$, $T$ denotes the temperature, and $\kappa$ denotes the thermal conductivity of the OFI, which is known to depend on the temperature. The boundary conditions are $T(0)=300$ and $T(1)=1900$. The resulting temperature measurements are denoted by $\mathbf{d}=(T_{i})_{i=1}^{51}$ and are assumed to be corrupted with zero-mean Gaussian noise with a standard deviation of $\sigma=2.5$ K. This results in an infinite-dimensional nonlinear inverse problem.
Under the classical approach, one would start by assigning a Gaussian measure to the thermal conductivity $\mu_{0}=\mathcal{N}(m,S)$, with $\log\kappa\sim\mu_{0}$ so that Bayes’s theorem can be used to derive the posterior. The measurement operator $R$ is given by the solution to eq. (5.2) at the measurement locations. Therefore the posterior $\mu^{\mathbf{d}}$ is given by the Radon-Nikodym derivative
@@
We demonstrate how this inverse problem can be solved by replacing the prior Gaussian measure with a BNN with the equivalent Mercer prior. For the prior knowledge, we need to impose the condition that $\kappa$ has at least one derivative, as $\partial\kappa/\partial T$ appears in eq. (5.2). Rather than defining the prior in the $\log$-space, we instead apply a softplus transformation, $\mathrm{softplus}(x)=\ln(1+e^{x})$, which helps prevent underflow. Putting this together, we select the prior $\mathrm{softplus}(\kappa)\sim\mathcal{N}(0,\Delta^{-2})$. We represent this prior as a BNN $\kappa_{\theta}$ with corresponding Mercer prior, with the eigenvalues and eigenfunctions being as discussed in Sec. 3.2. The BNN architecture in this example is a simple one layer network with $32$ neurons, where again we keep the first $100$ eigenpairs in the expansion. We find that the BNN does not need to be very expressive in order to produce prior samples which could reasonably represent $\kappa$, as evidenced in Fig. 15(a).
(a) Prior thermal conductivity samples
(b) Posterior thermal conductivity samples
@@
Before sampling from the posterior, we initialize the parameters at the maximum a posteriori estimate $\theta_{0}=\arg\max p(\theta|\mathbf{d})$. Each SGLD step uses $16$ randomly subsampled points in the domain to compute the $L^{2}$-inner products, $10$ randomly sampled eigenpairs, and a minibatch of $17$ temperature measurements, which is roughly $1/3$ of the data. We take $1$ million SGLD steps to characterize the posterior. The resulting BNN posterior is presented in Fig. 15, where we find that the posterior accurately reconstructs the latent thermal conductivity, although there is some discrepancy at higher temperatures. As this is an example with real experimental data, some amount of error is to be expected.
@@ 6 Conclusions
In this work, we introduced Mercer priors, a new class of priors for BNNs derived from the Mercer representation of a covariance kernel. The prior enforces a covariance structure on the parameters of the BNN such that in the output space the networks behave similarly to the corresponding GP with the chosen covariance. This allows for BNNs to inherit the interpretability of GP priors while retaining the scalability and flexibility of neural networks. One of the main advantages of the method is it works well for neural networks with relatively simple architectures and does not require much in the way of feature engineering. When it comes to selecting a model, the majority of the effort is spent in choosing an appropriate GP prior that we wish to approximate. A challenge that remains is the treatment of any hyperparameters appearing in the GP covariance, which was discussed in detail in Sec. 3.4.